OpenStudy (anonymous):

An archeologist finds an artifact that contains 44% of the carbon-14 that it would have had originally. What is the estimated age of the artifact? (The half-life of carbon-14 is 5,730 years.)

OpenStudy (mathstudent55):

Are you given a formula for calculating age using carbon-14 dating?

OpenStudy (anonymous):

no :(

OpenStudy (anonymous):

\[.44=\left(\frac{1}{2}\right)^\frac{t}{5730}\] solve for \(t\)

OpenStudy (anonymous):

in two steps \[\frac{t}{5730}=\frac{\ln(.44)}{\ln(.5)}\] \[t=\frac{5730\ln(.44)}{\ln(.5)}\] and a calculator

OpenStudy (anonymous):

thank you sooo much!! so is that the final answer?

OpenStudy (anonymous):

if you want a number, use a calculator

OpenStudy (anonymous):

to fine the estimated age

OpenStudy (anonymous):

looks like about 6787 http://www.wolframalpha.com/input/?i= \frac{5730\ln%28.44%29}{\ln%28.5%29}

OpenStudy (anonymous):

of the artifact

OpenStudy (anonymous):

i got t=6786.75

OpenStudy (anonymous):

maybe this one will work better http://www.wolframalpha.com/input/?i=5730*ln%28.44%29%2Fln%28.5%29

OpenStudy (anonymous):

thank you so much!!!

OpenStudy (anonymous):

yw